OpenStudy (anonymous):

Multiply and Simplify. 2mp • (–4p2) • (–2m2p)

OpenStudy (australopithecus):

add exponents of similar variables

OpenStudy (australopithecus):

hint in 2mp, m and p are to the power of 1. Also remember a (-)(+) = - and a (-)(-) = +

OpenStudy (anonymous):

2^1* (-4p^2) * (-2m2p) .....

OpenStudy (australopithecus):

2mp • (–4p2) • (–2m2p) = 2(-4)mp^(1)*p^(2) • (–2m2p) = -8mp^(3) • (–2m2p)

OpenStudy (australopithecus):

add exponents

OpenStudy (australopithecus):

and go one step at a time

OpenStudy (anonymous):

I Don't Get It -_-

OpenStudy (australopithecus):

what don't you get

OpenStudy (australopithecus):

2*2 = 4 2^(1)*2^(1) = 2^(2) = 2*2 = 4 2^(1) is the same thing as 2

OpenStudy (australopithecus):

when you see something like 3xpy realize it is just one number, we just dont know what x, p, or y is

OpenStudy (australopithecus):

when you see something like 3xpy realize it is just one number, we just dont know what x, p, or y is in terms of what they equal

OpenStudy (australopithecus):

i have no idea what you wrote

OpenStudy (australopithecus):

look it is vital you learn how to multiply variables. you can use this site to find the answer https://www.wolframalpha.com/ At least make an effort to understand what I'm telling you

OpenStudy (anonymous):

16m^3p^4 –16m^3p^4 8m^3p^4 –16I^2p^3

OpenStudy (australopithecus):

what do you think the answer is

OpenStudy (anonymous):

16m^3p^4

OpenStudy (australopithecus):

looks right

OpenStudy (anonymous):

Ok Thank You :)